OpenStudy (anonymous):

Identify the conic section from its equation. 1. (X-5)^2/4 + (y+3)^2/16 =1 2. (X-4)^2/36 - (y-2)^2/9 = 1 3. X^2+y^2+6x-16y+48=0 4. Y^2+6y-2x+13=0 5. (Y-2)^2/4 - (x+3)^2/9=0

OpenStudy (mathmale):

\[(x-5)^{2}/4 + (y+3)^{2}/16 = 1 => (x-5)^{2}/2^{2}+(y+3)^{2}/16 = 1\] matches the general equation of an ellipse:

OpenStudy (mathmale):

\[(x-h)^{2}/a ^{2}+(y-k)^{2}/b ^{2} = 1\]

OpenStudy (mathmale):

So the given equation #1 is the equation of an ellipse centered at (5,-3) with major axis 4 and minor axis 2. That is, the ellipse is a vertical one; the distance from one vertex to the other, vertically, is 2(4) = 8.